OpenStudy (anonymous):

Multiple choice :) The amount to which $1000 accumulates if invested for 2 years at 10% interest paid annually is? A. $1020.10 B. $1100 C. $1200 D. $1210

OpenStudy (wolfboy):

10% of a thousand is 100 right

OpenStudy (wolfboy):

100+1000=1100

OpenStudy (wolfboy):

10%1100 is 110

OpenStudy (anonymous):

1000*(1.1)^2=1210

OpenStudy (anonymous):

which means 1000*1.1=1100. 1100*1.1= 1210

OpenStudy (anonymous):

tyvm! i hate all money questions.. lol

OpenStudy (wolfboy):

Sorry I couldn't finish it for you my iPad froze up and got weird :(

OpenStudy (anonymous):

no prob :)